# Silent Guardian: Protecting Text from Malicious Exploitation by Large Language Models

### v3

@@ V-B2 Transfer Ability
V-B3 Time Discussion
@@
Text length. The text selected as the protection target is not an adjustable parameter. However, fortunately, due to the parallel computing nature of transformer models, the computational time for texts of different lengths remains relatively consistent. Similar to the experimental settings in Section V, we conducted 15 epochs of construction for texts in the Novel Dataset on the Vicuna-7B model. The average construction times in 40 tokens, 80 tokens and 120 tokens text were approximately 12.95 minutes, 13.15 minutes, and 15.18 minutes, respectively.
@@
Figure 6: The heatmap of the relationship between the replacement set size, batch size, and construction time. The masked areas indicate scenarios where, even after 100 epochs, the PSR did not reach 0.9 and the unit of the numbers in the figure is minutes.
@@
V-B4 Why Truncation Works
@@
Subsequently, we optimized the text to increase the probability of the model outputting the first token as the target token, iterating this process for ten rounds, while maintaining other experimental settings as outlined in Section V-A. The experimental results are depicted in Figure 7.
The results indicate that optimizing for the end token and “$\langle 0x0A\rangle$” as optimization targets was easily achieved, while other tokens were less sensitive to STP’s adjustments. Notably, some optimization targets initially exhibited higher values than the end token, gradually being overtaken by the end token across the epochs, highlighting the particularity of the end token. We refer to tokens sensitive to the optimization algorithm as “sensitive tokens” each corresponding to inherent properties of the model. For instance, the end token indicates a propensity to end dialogues, tokens ranking higher represent the model’s compliance with instructions, while tokens like “$\langle 0x0A\rangle$” correspond to certain biases generated during the model’s training. Exploring these biases may unveil deeper security implications.
Additionally, we discovered that optimizing the text to correspond to the token “$\langle 0x0A\rangle$” at index 13 also had a disruptive effect on the model’s generated results. We set the temperature to 0.7 and employed random sampling. To prevent excessively long text, we have set the maximum number of new tokens to 100. Some dialogue results are provided in the Appendix.
@@ VI-A Significance of This Work and Future Directions
Furthermore, TPE offers additional assurances for the rights of holders of high-quality data. Owners can employ targeted truncation optimizations on unregistered large models, thereby safeguarding their interests.
@@ VI-B STP Method
As shown in Figure 2, STP method ensures good concealment while allowing for rapid convergence. Importantly, unlike the GBDA method, STP does not require introducing a reference model to guarantee concealment, ensuring its applicability across various scenarios.
On the other hand, the STP method is an optimization of the GCG method, with a notable distinction. Unlike the GCG method, STP considers the concealment of adversarial text. Given that certain online models employ input detection mechanisms to prevent malicious usage [46], traditional methods such as adding prefixes or suffixes in prompts are susceptible to confusion detection. STP, in contrast, provides a less detectable avenue for evasion attacks, making it challenging to be identified.
@@ VI-C1 Different Models
For different models, the choice of the end token varies due to differences in tokenizers. The LLaMA model developed by Meta, for instance, uses the end token </s> On the other hand, models from the GPT [47] series typically include a special token, which incorporates the end token, when using their tokenizer. Taking the example of GPT-4’s cl100k-base [48] tokenizer, the corresponding dictionary index for the ending token is 100257. Therefore, when aiming to apply text truncation protection specifically for this model, this index should be used as the optimization target.
@@
VI-C2 Heightened Security Requirements
@@
To address the security requirements of explicitly defined text, optimization during the construction of TPE can target common prefixes and suffixes. Introducing coefficients to form a new loss function, optimization can be applied exclusively to the original text. For instance, when dealing with personal social media content, a defender may wish to conceal specific aspects of their lifestyle, and he can design corresponding prefixes and suffixes to construct TPE. The specific details are outlined in Algorithm 2.
VI-C3 Universal Model
@@
For models that share the same tokenizer but have different parameters, the approach aligns with the universal method outlined in the parentheses. This method enables simultaneous optimization for multiple models, thereby achieving enhanced model generality. The specific details are similar as Algorithm 2, except that the loss function calculation involves the sum across different models.
VI-C4 Other Truncation Method
@@
             $h_{i}=-\nabla_{v_{i}}\text{loss}$ $N_{i}=N(\mathcal{P}_{i})$ $S_{i}=Top-k(N_{i})$
      $\text{len}_{\text{part}}=\frac{B}{len(\mathcal{P})}$ for $b=1,...,B$ do
             $\lceil i=\frac{b}{\text{len}_{\text{part}}}\rceil$ $\tilde{\mathcal{P}}^{(b)}=\mathcal{P}$ $\tilde{\mathcal{P}}_{i}^{(b)}=Uniform(S_{i})$
@@ VII Conclusion
In this paper, we introduced the first text protection scheme, SG, designed to prevent malicious exploitation by LLM while safeguarding the privacy and copyright property of user-uploaded internet text through the concept of TPE. Our experimental outcomes demonstrated the effectiveness and concealment of the STP method across varied text lengths, types, and diverse models. Furthermore, TPE constructed using STP showed some level of transferability between models and robustness to prefixes and suffixes. We also explored the time required to construct TPE using STP, finding that with appropriately chosen parameters, the construction time for individual protection is remarkably short. Additionally, we delved into optimizing specific token-associated model properties, which we believe can inspire future investigations into LLM characteristics. We aim to deploy SG in real-world scenarios to address the escalating concerns regarding LLM security and believe that it will find broader applications in the future.

### v4

@@ V-B2 Transferability
In the Novel dataset, experimental results show that transferability improves with the increase in text length. However, in the Vicuna dataset, we observed that the TPE optimized for different models exhibited weaker transferability between models. To address this issue, we aggregated the loss functions of different models into a new loss function to construct TPE, hoping to enhance transferability to some extent.
Specifically, for three models $\text{model}_{1}$, $\text{model}_{2}$, and $\text{model}_{3}$, we defined the new loss function as $(\text{loss}_{\text{model}_{1}}+\text{loss}_{\text{model}_{2}})/2$, and used this loss function to construct TPE, which is used to attack $\text{model}_{3}$. The specific experimental results are shown in Table VIII. Here, AVE represents the average of the results in the first two rows of the table, and AGG represents the aggregating of the other two models.
The experimental results show that the TPE constructed using the aforementioned method exhibits enhanced transferability. Compared to the original method, the average PSR values across various models in the Vicuna dataset have increased by 0.11. Furthermore, from the above experimental results, it can be reasonably inferred that aggregating the loss functions of more models to construct TPE should lead to even better transferability.
@@
TABLE VIII: Enhancement of TPE Transferability on the Vicuna Dataset
Model Metrics Vicuna Dataset
Guanaco PSR 0.56 0.53 0.51 0.67 0.27 0.22 0.70 0.45 0.60
[…]
LLaMA PSR 0.53 0.46 0.41 0.65 0.22 0.39 0.44 0.24 0.47
[…]
Vicuna PSR 0.97 0.95 0.88 1.00 0.63 0.79 0.99 0.79 0.88
[…]
V-B3 The Relationship between the Time and Hyperparameters for Constructing a Single TPE
@@
Text length. The text selected as the protection target is not an adjustable parameter. However, fortunately, due to the parallel computing nature of transformer models, the computational time for texts of different lengths remains relatively consistent. Similar to the experimental settings in Section V, we conducted 15 epochs of construction for texts in the Novel Dataset on the Vicuna-7B model. The average construction times in 40 tokens, 80 tokens, and 120 tokens text were approximately 12.95 minutes, 13.15 minutes, and 15.18 minutes, respectively.
@@
Figure 6: The heatmap of the relationship between the replacement set size, batch size, and construction time. The masked areas indicate scenarios where, even after 100 epochs, the PSR did not reach 0.9, and the unit of the numbers in the figure is minutes.
@@
V-B4 Discussion on the Relationship between TPE Construction Time and Text Length
In this section, we will discuss the relationship between the time required for TPE construction and the length of the text. Specifically, we conducted experiments on texts of lengths 40, 80, 120, 160, 200, 240, 280, and 320 tokens from the Novel dataset. The experimental parameters were set as $|S_{i}|=5$, $|N_{i}|=10$, $batch\ size=128$, $T=15$. The experimental results are shown in Figure 7. All experiments in this section were completed using an NVIDIA RTX A6000 GPU.
Figure 7: The relationship between TPE construction time and text length.
The experimental results indicate that the time cost is approximately linearly related to the length of the text. Furthermore, as discussed in Section V-B3 of the manuscript, adjusting the parameter settings appropriately can further reduce the time cost. Additionally, our text protection method can be applied in offline scenarios without the need for real-time computation, making its application in real-world scenarios feasible.
V-B5 Why Truncation Works
@@
Subsequently, we optimized the text to increase the probability of the model outputting the first token as the target token, iterating this process for ten rounds while maintaining other experimental settings as outlined in Section V-A. The experimental results are depicted in Figure 8.
The results indicate that optimizing for the end token and “$\langle 0x0A\rangle$” as optimization targets were easily achieved, while other tokens were less sensitive to STP’s adjustments. Notably, some optimization targets initially exhibited higher values than the end token, gradually being overtaken by the end token across the epochs, highlighting the particularity of the end token. We refer to tokens sensitive to the optimization algorithm as “sensitive tokens,” each corresponding to inherent properties of the model. For instance, the end token indicates a propensity to end dialogues, tokens ranking higher represent the model’s compliance with instructions, while tokens like “$\langle 0x0A\rangle$” correspond to certain biases generated during the model’s training. Exploring these biases may unveil deeper security implications.
@@ VI-B STP Method
As shown in Figure 2, the STP method ensures good concealment while allowing for rapid convergence. Importantly, unlike the GBDA method, STP does not require introducing a reference model to guarantee concealment, ensuring its applicability across various scenarios.
On the other hand, the STP method is an optimization of the GCG method, with a notable distinction. Unlike the GCG method, STP considers the concealment of adversarial text. Given that certain online models employ input detection mechanisms to prevent malicious usage [44], traditional methods such as adding prefixes or suffixes in prompts are susceptible to perplexity detection. STP, in contrast, provides a less detectable avenue for jailbreak attacks, making it challenging to be identified.
@@ VI-C1 Different Models
The specified end token is not the same for different models due to tokenizer differences. For example, the LLaMA model developed by Meta uses </s> as the end token, which corresponds to a dictionary index of 2, while the end token in the cl100k-base [45] tokenizer of the GPT-4 model corresponds to a dictionary index of 100257. Therefore, different indexes should be used as the optimization target when constructing TPEs for different models.
@@
VI-C2 Discussion on the Robustness of TPE
In this section, we will discuss the robustness of TPE. Specifically, we selected the experimental results from Section V-B1, concerning 120-token length texts from the Novel dataset optimized on the Vicuna model and subjected them to random word deletion, random word addition, and random synonym replacement operations [46]. The randomization ratio ranged from 5% to 20%. The experimental results are presented in Table IX, where RI denotes random word insertion, RD denotes random word deletion, and SR denotes random synonym replacement, with subscripts indicating the random operation ratio.
TABLE IX: Robustness Performance of TPE
Metric Operations None $\text{RI}_{5\%}$ $\text{RI}_{10\%}$ $\text{RI}_{15\%}$ $\text{RI}_{20\%}$
PSR 0.80 0.51 0.39 0.33 0.27
Metric Operations None $\text{RD}_{5\%}$ $\text{RD}_{10\%}$ $\text{RD}_{15\%}$ $\text{RD}_{20\%}$
PSR 0.80 0.48 0.33 0.23 0.15
Metric Operations None $\text{SR}_{5\%}$ $\text{SR}_{10\%}$ $\text{SR}_{15\%}$ $\text{SR}_{20\%}$
PSR 0.80 0.51 0.37 0.29 0.25
The results indicate that the effectiveness of TPE protection deteriorates as the perturbation ratio increases. However, excessively high ratios of perturbation can cause significant loss to the original text.
VI-C3 Heightened Security Requirements
@@
For text with clearly defined security requirements, the TPE construction can be optimized by focusing on the prefixes and suffixes related to the specific topic. A new loss function can be constructed by introducing coefficients, and optimization should be applied only to the original text portion. For instance, when dealing with personal social media content, a defender may wish to conceal specific aspects of their lifestyle, and he can design corresponding prefixes and suffixes to construct TPE. The specific details are outlined in Algorithm 2.
In this section, we prepared five rewrite-related prefixes, denoted as $\text{Prefix}_{1}^{\prime}$, $\text{Prefix}_{2}^{\prime}$, $\text{Prefix}_{3}^{\prime}$, $\text{Prefix}_{4}^{\prime}$, and $\text{Prefix}_{5}^{\prime}$, along with 100 rewrite-related prefixes generated by ChatGPT-3.5. We defined the loss function as $(\sum_{i=1}^{5}\text{loss}_{\text{Prefix}_{i}^{\prime}})/5$. The experimental parameters were set as $|S_{i}|=5$, $|N_{i}|=10$, $batch\ size=128$, and $T=15$. The experiments were conducted on the Vicuna model using texts with a length of 40 tokens from the Novel dataset. The experimental results are shown in Table X, where $\text{PSR}^{\dagger}$ represents the average results after transferring to the 100 rewrite-related prefixes.
The experimental results show that under the theme of rewriting, TPE maintains a high PSR value even after transferring to unoptimized prefixes, with an average PSR value of 0.81 after transfer.
The specific prefixes are provided in the Appendix.
TABLE X: Transferability of Prefix Optimization
Metrics Prefix $\text{Prefix}_{1}^{\prime}$ $\text{Prefix}_{2}^{\prime}$ $\text{Prefix}_{3}^{\prime}$ $\text{Prefix}_{4}^{\prime}$ $\text{Prefix}_{5}^{\prime}$
PSR 0.88 0.91 0.89 0.90 0.91
$\text{PSR}^{\dagger}$ 0.81
VI-C4 Discussion on TPE Text Quality
In the aforementioned STP method, we used n-nearest neighbors of tokens in the embedding layer to construct semantically similar candidate sets. In fact, the construction of semantically similar candidate sets in STP is a replaceable module, and different methods can be used to construct semantically similar candidate sets to improve the quality of TPE texts. In this section, we mainly discuss the impact of using BERT to construct [27] these candidate sets on the quality of TPE texts.
Specifically, we used the BERT-base-uncased model to construct semantically similar sets. Compared to the original method, this method improved the semantic fluency of the text without compromising optimization efficiency, although there remains a large semantic gap compared to the original text. Some experimental results on the Vicuna model are presented in Table XI.
TABLE XI: Examples of TPE on Vicuna with BERT
Original Text Protected Text
Can you help me write a resignation letter to my current employer, while leaving on good terms and expressing gratitude for the opportunities provided? Can you exc me , a resignation letter to my previous employ es, for leaving on our terms and expressing gratitude to the opportunities provided?
Draft an apology email to a customer who experienced a delay in their order, and provide reassurance that the issue has been resolved. Draft an apology email by a student which experienced a change in their order, and seek anassurance that his issue previously remained cleared !
Write a compelling product launch announcement email to inform our customers of our new software solution. Write a jelling product launch announ cy email to inform my staff of our current hardware deployment ?
Write a captivating movie review for a recently released science fiction film, discussing its plot, characters, and special effects. Write a captivating feature review for a newly released action comedy film, learning about unusual situations, characters background & software engineering.
VI-C5 Universal Model
@@
For models that share the same tokenizer but have different parameters, the approach aligns with the universal method outlined in the GCG method[19]. This method enables simultaneous optimization for multiple models, thereby achieving enhanced TPE generality. The specific details are similar to Algorithm 2, except that the loss function calculation involves the sum across different models.
VI-C6 Other Truncation Method
@@
             $h_{i}=-\nabla_{v_{i}}\text{loss}$
             $N_{i}=N(\mathcal{P}_{i})$
             $S_{i}=Top-k(N_{i})$
      $\text{len}_{\text{part}}=\frac{B}{len(\mathcal{P})}$
       for $b=1,...,B$ do
             if $B>\text{len}(\mathcal{P})$ then
                  $i=\lceil\frac{b}{\text{len}_{\text{part}}}\rceil$
             else
                   repeat
                         $i=\text{random}(1,len(\mathcal{P}))$
                  until $i\neq\text{previous }i$
             $\tilde{\mathcal{P}}^{(b)}=\mathcal{P}$
             $\tilde{\mathcal{P}}_{i}^{(b)}=Uniform(S_{i})$
@@ VII Conclusion
In this paper, we introduce the first text protection mechanism, SG, and the concept of TPE to address the problem of malicious exploitation of texts by LLMs. Meanwhile, we introduce the first method for constructing TPE, which is called STP. Our experimental outcomes demonstrated the effectiveness and concealment of the STP method across varied text lengths, types, and diverse models. Furthermore, TPE constructed using STP showed some level of transferability and robustness. Additionally, we explore the time overhead of constructing TPE using the STP method, which, under appropriately chosen parameters, incurs very short construction times for individual texts and exhibits a nearly linear relationship with text length. Additionally, we delved into optimizing specific token-associated model properties, which we believe can inspire future investigations into LLM characteristics. We aim to deploy SG in real-world scenarios to address the escalating concerns regarding LLM security and believe that it will find broader applications in the future.